# Graph Transformers Dream of Electric Flow

### v1

@@ 3.3.1 Computing The Principal Square Root $\sqrt{{\mathcal{L}}^{\dagger}}$
One can verify the above by noticing that ${\left\|{\mathcal{M}}_{i}-{\mathcal{M}}_{j}\right\|}_{2}^{2}={\mathcal{M}}_{i}^{\top}{\mathcal{M}}_{i}+{\mathcal{M}}_{j}^{\top}{\mathcal{M}}_{j}-2{\mathcal{M}}_{i}^{\top}{\mathcal{M}}_{j}={\left[{\mathcal{L}}\right]}_{ii}+{\left[{\mathcal{L}}\right]}_{jj}-2{\mathcal{L}}_{ij}$. In some settings, it is more natural to use the rows of ${\mathcal{M}}$ to embed node position, instead of directly using ${\mathcal{R}}$: By assigning an embedding vector $w_{i}:={\mathcal{M}}_{i}$ to vertex $i$, the Euclidean distance between $w_{i}$ and $w_{j}$ equals the resistance distance.
The matrix ${\mathcal{M}}$ is under-determined, and for any $m>n$, there are infinitely many choices of ${\mathcal{M}}$ that satisfy ${\mathcal{M}}{\mathcal{M}}^{\top}={\mathcal{L}}^{\dagger}$. Fact 1 applies to any such ${\mathcal{M}}$. Velingker et al. (2024) uses the rows of ${\mathcal{M}}={\mathcal{L}}^{\dagger}B$ for resistive embedding. Under this choice, ${\mathcal{M}}_{i}$ has dimension $d$, which is the number of edges and can be quite large. To deal with this, Velingker et al. (2024) additionally performs a dimension-reduction step using Johnson Lidenstrauss.
@@
Lemma 2 (Principal Square Root $\sqrt{{\mathcal{L}}^{\dagger}}$).
Consider the setup in Section 3.1. Assume that $\psi_{1}...\psi_{k}\in\mathbb{R}^{n}$ satisfy $\langle{\psi_{i},\vec{1}}\rangle=0$. For any $L$-layer Transformer (2.3), there exists a configuration of weights $W^{V},W^{Q},W^{K},W^{R}$ such that the following holds: For any graph with maximum Laplacian eigenvalue less than $\lambda_{\max}$ and minimum non-trivial Laplacian eigenvalue greater than $\lambda_{\min}$, and for all $i=1...k$:
@@ 3.3.2 Computing the heat kernel: $\exp{\left(-s{\mathcal{L}}\right)}$
Finally, we present a result on learning heat kernels. The heat kernel has connections to random walks and diffusion maps Coifman and Lafon (2006). It plays a central role in semi-supervised learning on graphs (Xu et al., 2020); it is also used for positional encoding (Choromanski et al., 2022). We note that in some settings, the heat kernel excludes the nullspace of ${\mathcal{L}}$ and is instead defined as $e^{-s{\mathcal{L}}}-\vec{1}\vec{1}^{\top}/n$.
Lemma 3 (Heat Kernel $e^{-s{\mathcal{L}}}$).
Consider the setup in Section 3.1. Assume that $\psi_{1}...\psi_{k}\in\mathbb{R}^{n}$ satisfy $\langle{\psi_{i},\vec{1}}\rangle=0$. Let $s>0$ be an arbitrary temperature parameter. There exists a configuration of weights for the $L$-layer Transformer (2.3) such that the following holds: for any input graph whose Laplacian ${\mathcal{L}}$ satisfies $8s\lambda_{\max}\leq L$, and for all $i=1...k$,
@@ 3.4 Implementing Multiplicative Polynomial Expansion
Lemma 4.
Let $\delta>0$. Let $Z_{0}^{\top}:=\begin{bmatrix}(\hat{I}_{n\times n}-\delta{\mathcal{L}}),&I_{n\times n},&\delta I_{n\times n}\end{bmatrix}$. Then there exist a choice of $W^{V},W^{Q},W^{K},W^{R}$ for a $L$-layer Transformer (2.3) such that for any graph with Laplacian with smallest non-trivial eigenvalue $\lambda_{\min}$ and largest eigenvalue $\lambda_{\max}\leq 1/\delta$,
@@
Lemma 5 (Fast Heat Kernel).
@@ 3.5 Experiments
In Figure 1, we experimentally verify that the Transformer is capable of learning to solve the three objectives presented in Lemmas 1, 2 and 3. The setup is as described in Section 3.1, with the Transformer described in (2.3) with $k=n$, but with the following important difference: the output of each layer is additionally normalized as follows: ${\left[Z_{l}\right]}_{1...n}\leftarrow{\left[Z_{l}\right]}_{1...n}/\|{\left[Z_{l}\right]}_{1...n}\|_{F}$, ${\left[Z_{l}\right]}_{n+1...n+2k}\leftarrow{\left[Z_{l}\right]}_{n+1...n+2k}/\|{\left[Z_{l}\right]}_{n+1...n+2k}\|_{F}$, where $\|\|_{F}$ is the Frobenius norm. Without this normalization, training the Transformer becomes very difficult beyond 5 layers.
@@ 4 Transformers can Implement Subspace Iteration to Compute Eigenvectors
We present a Transformer construction for finding the eigenvectors of the graph Laplacian ${\mathcal{L}}$. The eigenvector for the smallest non-trivial eigenvalue, for example, has been applied with great success in graph segmentation (Shi and Malik, 1997) and clustering (Bühler and Hein, 2009). Additionally, the Laplacian eigenvectors are also commonly used for positional encoding (see Section 1.2). Our construction is based on the subspace iteration algorithm (Algorithm 1), aka block power method—see e.g., (Bentbib and Kanber, 2015). The output $\Phi$ of Algorithm 1 converges to the top-$k$ eigenvectors of ${\mathcal{L}}$. We show in Corollary 7 that a modified construction can also find the bottom-$k$ eigenvectors of ${\mathcal{L}}$.
@@
$\Phi_{0}\in\mathbb{R}^{n\times k}$ has full column rank .
@@
     $\hat{\Phi}\leftarrow{\mathcal{L}}\Phi$
     $\Phi\leftarrow QR(\Phi)$triangleright QR($\Phi$) returns the Q matrix in the QR decomposition of $\Phi$.
@@
For the purposes of this section, we consider a variant of the Transformer defined in (2.3).
@@
${Z}_{l+1}=\text{normalize}(\hat{Z}_{l+1}),$ (4.1)
where $\text{normalize}(Z)$ applies row-wise normalization: ${\left[\text{normalize}(Z)\right]}_{i}\leftarrow{\left[Z\right]}_{i}/\|{\left[Z\right]}_{i}\|_{2}$ for $i=d+1...d+k$. We use $B_{l}\in\mathbb{R}^{n\times d}$ and $\Phi_{l}\in\mathbb{R}^{n\times k}$ to denote refer to specific columns of $Z_{l}^{\top}$, defined as $Z_{l}^{\top}=:\begin{bmatrix}B_{l}^{T},&\Phi_{l}^{\top}\end{bmatrix}$. The notation is chosen as $\Phi_{l}$ corresponds to $\Phi$ in Algorithm 1. We initialize $B_{0}=B$ and let $\Phi_{0}$ be any column-orthogonal matrix.
Lemma 6 (Subspace Iteration for Finding Top $k$ Eigenvectors).
Consider the Transformer defined in (4.1). There exists a choice of $W^{V},W^{Q},W^{K},W^{R}$ such that $k+1$ layers of the Transformer implements one iteration of Algorithm 1. Consequently, the output $\Phi_{L}$ of a $L$-layer Transformer approximates the top-$k$ eigenvectors of ${\mathcal{L}}$ to the same accuracy as $L/(k+1)$ steps of Algorithm 1.
Discussion. We bound the Transformer’s error by the error of Algorithm 1, but do not provide an explicit convergence rate. This omission is because the convergence rate of Algorithm 1 itself is difficult to characterize, and has a complicated dependence on pairwise spectral gaps. The high-level proof idea is to use self-attention to orthogonalize the columns of $\Phi_{l}$ (see Lemma 8). We defer the proof of Lemma 6 to Appendix 9. We experimentally validate Lemma 6 in Section 4.1.
@@
In graph applications, one is often interested in the smallest-$k$ eigenvectors of ${\mathcal{L}}$:
Corollary 7 (Subspace Iteration for Finding Bottom $k$ Eigenvectors).
Consider the same setup as Lemma 6. Let $\mu>0$ be some constant. There exists a construction for a $L$-layer Transformer (similar to Lemma 6), which implements Algorithm 1 with ${\mathcal{L}}$ replaced by $\mu I-{\mathcal{L}}$, and $\Phi_{L}$ approximates the smallest $k$ eigenvectors of ${\mathcal{L}}$ if $\lambda_{\max}({\mathcal{L}})\leq\mu$.
We provide a short proof in Appendix 9. The key idea is that a minor modification of the contruction in Lemma 6 will instead compzute the bottom $k$ eigenvectors. Alternative to the construction in Corollary 7, one can also first compute ${\mathcal{L}}^{\dagger}$ (via Lemma 4), and then perform subspace iteration for ${\mathcal{L}}^{\dagger}$.
@@ 4.1 Experiments for Lemma 6
We verify Lemma 6 and Corollary 7 experimentally by evaluating the ability of the Transformer (2.3) to learn top-$k$ and bottom-$k$ eigenvectors. As in Section 3.5, we consider two kinds of random graphs with $n=10$ nodes: fully connected ($d=45$ edges) and CSL ($d=20$ edges); each edge is has a randomly sampled resistance; see Appendix 11 for details. For a graph ${\mathcal{G}}$ with Laplacian ${\mathcal{L}}$, let $\lambda_{1}\leq\lambda_{2}\leq...\lambda_{10}$ denote its eigenvalues. Let $v_{1},\dots,v_{10}$ denote its eigenvectors. $\lambda_{1}$ is always 00 and $v_{1}$ is always $\vec{1}/\sqrt{n}$.
The Transformer architecture is as defined in Section 4.1, with $k=n$. We increase the dimension of $\Phi_{l}$ to $(2n)\times n$, and thus the dimension of $Z_{l}$ to $(d+n+n)\times n$. We read out the last $n$ rows of $Z_{L}$ as output. The purpose of increasing the dimension is to make the architecture identical to the one used in the experiments in Section 3.5; the construction in Lemmas 6 and Corollary 7 extend to this setting by setting appropriate parameters to 00. In addition, we also normalize ${\left[Z_{l}\right]}_{1...d}$ each layer by its Frobenius norm, i.e., ${\left[Z_{l}\right]}_{1...d}\leftarrow{\left[Z_{l}\right]}_{1...d}/\|{\left[Z_{l}\right]}_{1...d}\|_{F}$ (the proof of Lemma 6 still holds as subspace iteration is scaling-invariant). The input to the Transformer is $Z_{0}^{\top}=\begin{bmatrix}B,&\Phi_{0}\end{bmatrix}$, and we make $\Phi_{0}\in\mathbb{R}^{2n\times n}$ a trainable parameter along with $W^{V},W^{Q},W^{K},W^{R}$. We define $\text{loss}_{i}:=\mathbb{E}{\left[\min\left\{\|\phi_{L,i}-v_{i}\|_{2}^{2},\|\phi_{L,i}+v_{i}\|_{2}^{2}\right\}\right]}$, where $\phi_{L,i}$ is the $i^{th}$ row of $\Phi_{L}$ and expectation is taken with respect to randomness in sampling the graph; the $\min$ is required because eigenvectors are invariant to sign-flips. We train and evaluate the Transformer on two kinds of losses: $\text{loss}_{1-5}:=\frac{1}{5}\sum_{i=1}^{5}\text{loss}_{i}$ and $\text{loss}_{1-10}:=\frac{1}{10}\sum_{i=1}^{10}\text{loss}_{i}$. We plot the results in Figure 2, and summarize our findings below:
@@ 5 Parameter-Efficient Implementation
As explained in Section 3.1, the sizes of the $W^{Q},W^{K},W^{V},W^{R}$ matrices can scale with $O(n^{4})$ in the worst case, where $n$ is the number of nodes. This makes the memory requirement prohibitive. To alleviate this, we present below in (5.1) a more efficient Transformer implementation than the one defined in (2.3). (5.1) is strictly more constrained than (2.3), but it is still expressive enough to implement all the previous constructions. For each layer $l$, let $\alpha^{V}_{l}$, $\alpha^{Q}_{l}$, $\alpha^{K}_{l}$, $\alpha^{R}_{l}$ be scalar weight parameters, and let $W_{l}^{V,\Phi}$, $W_{l}^{Q,\Phi}$, $W_{l}^{K,\Phi}$, $W_{l}^{R,\Phi}\in\mathbb{R}^{2k\times 2k}$. Let $B_{l},\Phi_{l}$ evolve as
$B_{l+1}^{\top}=(1+\alpha^{R}_{l})B_{l}^{\top}+\alpha^{V}_{l}B_{l}^{\top}{\left(\alpha^{Q}_{l}\alpha^{K}_{l}B_{l}B_{l}^{\top}+\Phi_{l}{W_{l}^{Q,\Phi}}^{\top}W_{l}^{K,\Phi}\Phi_{l}^{\top}\right)}$ (5.1)
@@
with initialization $B_{0}:=B$, and $\Phi_{0}=[\Psi,0_{k\times n}]$. We verify that (5.1) matches (2.3), with $Z_{l}=\begin{bmatrix}B_{l}^{\top},\Phi_{l}^{\top}\end{bmatrix}$; the difference is that certain blocks of $W^{V}_{l}$, $W^{Q}_{l}$, $W^{K}_{l}$, $W^{R}_{l}$ constrained to be zero or some scaling of identity. Nonetheless, we verify that the constructions in Lemmas 1, 2, 3 and 6 can all be realized within the constrained dynamics (5.1). We prove this in Lemma 9 in Appendix 10.
Besides the reduced parameter size, we highlight two additional advantages of (5.1):
@@
    Let ${\sf TF}^{B}_{l}({\color[rgb]{0,0,1}B},{\color[rgb]{0,0,1}\Phi})$ (resp ${\sf TF}^{\Phi}_{l}({\color[rgb]{0,0,1}B},{\color[rgb]{0,0,1}\Phi}))$ be defined as the value of $B_{l}^{\top}$ (resp $\Phi_{l}^{\top}$) when initialized at $B_{0}={\color[rgb]{0,0,1}B}$ and $\Phi_{0}={\color[rgb]{0,0,1}\Phi}$ under the dynamics (5.1). Let $U\in\mathbb{R}^{d\times d}$ be some permutation matrix over edge indices. Then

### v2

@@ 3.3.1 Computing The Principal Square Root $\sqrt{{\mathcal{L}}^{\dagger}}$
One can verify the above by noticing that ${\left\|{\mathcal{M}}_{i}-{\mathcal{M}}_{j}\right\|}_{2}^{2}={\mathcal{M}}_{i}^{\top}{\mathcal{M}}_{i}+{\mathcal{M}}_{j}^{\top}{\mathcal{M}}_{j}-2{\mathcal{M}}_{i}^{\top}{\mathcal{M}}_{j}={\left[{\mathcal{L}}\right]}_{ii}+{\left[{\mathcal{L}}\right]}_{jj}-2{\mathcal{L}}_{ij}$. In some settings, it is more natural to use the rows of ${\mathcal{M}}$ to embed node position, instead of directly using ${\mathcal{R}}$: By assigning an embedding vector $w_{i}:={\mathcal{M}}_{i}$ to vertex $i$, the Euclidean distance between $w_{i}$ and $w_{j}$ equals the resistance distance. The matrix ${\mathcal{M}}$ is under-determined, and for any $m>n$, there are infinitely many choices of ${\mathcal{M}}$ that satisfy ${\mathcal{M}}{\mathcal{M}}^{\top}={\mathcal{L}}^{\dagger}$. Fact 1 applies to any such ${\mathcal{M}}$. Velingker et al. (2024) uses the rows of ${\mathcal{M}}={\mathcal{L}}^{\dagger}B$ for resistive embedding. Under this choice, ${\mathcal{M}}_{i}$ has dimension $d=|{\mathcal{E}}|$, which can be quite large. To deal with this, they additionally performs a dimension-reduction step using Johnson Lidenstrauss.
@@
Lemma 2 (Principal Square Root $\sqrt{{\mathcal{L}}^{\dagger}}$)
Consider the setup in Section 3.1. Assume that $\psi_{1}...\psi_{k}\in\mathbb{R}^{n}$ satisfy $\langle{\psi_{i},\vec{1}}\rangle=0$. For any $L$-layer Transformer (3), there exists a configuration of weights $W^{V},W^{Q},W^{K},W^{R}$ such that the following holds: For any graph with maximum Laplacian eigenvalue less than $\lambda_{\max}$ and minimum non-trivial Laplacian eigenvalue greater than $\lambda_{\min}$, and for all $i=1...k$:
@@ 3.3.2 Computing the heat kernel: $\exp{\left(-s{\mathcal{L}}\right)}$
Finally, we present a result on learning heat kernels. The heat kernel has connections to random walks and diffusion maps Coifman & Lafon (2006). It plays a central role in semi-supervised learning on graphs (Xu et al., 2020); it is also used for positional encoding (Choromanski et al., 2022). Note that sometimes, the heat kernel excludes the nullspace of ${\mathcal{L}}$ and is defined as $e^{-s{\mathcal{L}}}-\vec{1}\vec{1}^{\top}/n$.
Lemma 3 (Heat Kernel $e^{-s{\mathcal{L}}}$)
Consider the setup in Section 3.1. Assume that $\psi_{1}...\psi_{k}\in\mathbb{R}^{n}$ satisfy $\langle{\psi_{i},\vec{1}}\rangle=0$. Let $s>0$ be an arbitrary temperature parameter. There exists a configuration of weights for the $L$-layer Transformer (3) such that the following holds: for any input graph whose Laplacian ${\mathcal{L}}$ satisfies $8s\lambda_{\max}\leq L$, and for all $i=1...k$,
@@ 3.4 Implementing Multiplicative Polynomial Expansion
Lemma 4
Let $\delta>0$. Let $Z_{0}^{\top}:=\begin{bmatrix}(\hat{I}_{n\times n}-\delta{\mathcal{L}}),&I_{n\times n},&\delta I_{n\times n}\end{bmatrix}$. Then there exist a choice of $W^{V},W^{Q},W^{K},W^{R}$ for a $L$-layer Transformer (3) such that for any graph with Laplacian with smallest non-trivial eigenvalue $\lambda_{\min}$ and largest eigenvalue $\lambda_{\max}\leq 1/\delta$,
@@
Lemma 5 (Fast Heat Kernel)
@@ 3.5 Experiments
In Figure 1, we experimentally verify that the Transformer is capable of learning to solve the three objectives presented in Lemmas 1, 2 and 3. The setup is as described in Section 3.1, with the Transformer described in (3) with $k=n$, but with the following important difference: the output of each layer is additionally normalized as follows: ${\left[Z_{l}\right]}_{1...n}\leftarrow{\left[Z_{l}\right]}_{1...n}/\|{\left[Z_{l}\right]}_{1...n}\|_{F}$, ${\left[Z_{l}\right]}_{n+1...n+2k}\leftarrow{\left[Z_{l}\right]}_{n+1...n+2k}/\|{\left[Z_{l}\right]}_{n+1...n+2k}\|_{F}$, where $\|\|_{F}$ is the Frobenius norm. Without this normalization, training the Transformer becomes very difficult beyond 5 layers.
@@ 4 Transformers can implement Subspace Iteration to compute Eigenvectors
We present a Transformer construction for computing the eigenvectors of the graph Laplacian ${\mathcal{L}}$. The Laplacian eigenvectors are commonly used for positional encoding (see Section 1.2). In particular, the eigenvector for the smallest non-trivial eigenvalue has been applied with great success in graph segmentation (Shi & Malik, 1997) and clustering (Bühler & Hein, 2009). Our construction is based on the subspace iteration algorithm (Algorithm 1), aka block power method—see (Bentbib & Kanber, 2015). The output $\Phi$ of Algorithm 1 converges to the top-$k$ eigenvectors of ${\mathcal{L}}$. In Corollary 7, we present a modified construction that instead computes the bottom-$k$ eigenvectors of ${\mathcal{L}}$.
@@
$\Phi_{0}\in\mathbb{R}^{n\times k}$ has full column rank
@@
     $\hat{\Phi}_{i+1}\leftarrow{\mathcal{L}}\Phi_{i}$
     $\Phi_{i+1}\leftarrow QR(\hat{\Phi}_{i+1})$triangleright QR($\Phi$) returns the Q matrix in the QR decomposition of $\Phi$.
@@
For the purposes of this section, we consider a variant of the Transformer defined in (3).
@@
${Z}_{l+1}=\text{normalize}(\hat{Z}_{l+1}),$ (6)
where $\text{normalize}(Z)$ applies row-wise normalization: ${\left[\text{normalize}(Z)\right]}_{i}\leftarrow{\left[Z\right]}_{i}/\|{\left[Z\right]}_{i}\|_{2}$ for $i=d+1...d+k$. We use $B_{l}\in\mathbb{R}^{n\times d}$ and $\Phi_{l}\in\mathbb{R}^{n\times k}$ to denote refer to specific columns of $Z_{l}^{\top}$, defined as $Z_{l}^{\top}=:\begin{bmatrix}B_{l}^{T},&\Phi_{l}^{\top}\end{bmatrix}$. The notation is chosen as $\Phi_{l}$ in $Z_{l}$ in (6) corresponds to $\Phi_{l}$ in Algorithm 1. We initialize $B_{0}=B$ and let $\Phi_{0}$ be some arbitrary matrix with full column rank. $\Phi_{1}$ becomes orthogonal after the first QR step of Algorithm 1.
Lemma 6 (Subspace Iteration for Finding Top $k$ Eigenvectors)
Consider the Transformer defined in (6). There exists a choice of $W^{V},W^{Q},W^{K},W^{R}$ such that $k+1$ layers of the Transformer implements one iteration of Algorithm 1. Consequently, the output $\Phi_{L}$ of a $L$-layer Transformer approximates the top-$k$ eigenvectors of ${\mathcal{L}}$ to the same accuracy as $L/(k+1)$ steps of Algorithm 1.
Discussion. We bound the Transformer’s error by the error of Algorithm 1, but do not provide an explicit convergence rate. This omission is because the convergence rate of Algorithm 1 itself is difficult to characterize, and has a complicated dependence on pairwise spectral gaps of $\Phi_{0}$. The high-level proof idea is to use self-attention to orthogonalize the columns of $\Phi_{l}$ (Lemma 8. We defer the proof of Lemma 6 to Appendix 11. We experimentally validate Lemma 6 in Section 4.1.
@@
In graph applications, one is often interested in the bottom-$k$ eigenvectors of ${\mathcal{L}}$. The following corollary shows that a minor modification of Lemma 6 will instead compute the bottom $k$ eigenvectors.
Corollary 7 (Subspace Iteration for Finding Bottom $k$ Eigenvectors)
Consider the same setup as Lemma 6. Let $\mu>0$ be some constant. There exists a construction for a $L$-layer Transformer (similar to Lemma 6), which implements Algorithm 1 with ${\mathcal{L}}$ replaced by $\mu I-{\mathcal{L}}$, and $\Phi_{L}$ approximates the bottom $k$ eigenvectors of ${\mathcal{L}}$ if $\lambda_{\max}({\mathcal{L}})\leq\mu$.
We provide a short proof in Appendix 11. The key idea is that a minor modification of the construction in Lemma 6 will instead compute the bottom $k$ eigenvectors. Alternative to the construction in Corollary 7, one can also first compute ${\mathcal{L}}^{\dagger}$ (via Lemma 4), followed by subspace iteration for ${\mathcal{L}}^{\dagger}$.
@@ 4.1 Experiments for Lemma 6
We verify Lemma 6 and Corollary 7 experimentally by evaluating the ability of the Transformer (3) to learn top-$k$ and bottom-$k$ eigenvectors. As in Section 3.5, we consider two kinds of random graphs with $n=10$ nodes: fully connected ($d=45$ edges) and CSL ($d=20$ edges); each edge is has a randomly sampled resistance; see Appendix 13 for details. For a graph ${\mathcal{G}}$ with Laplacian ${\mathcal{L}}$, let $\lambda_{1}\leq\lambda_{2}\leq...\lambda_{10}$ denote its eigenvalues. Let $v_{1},\dots,v_{10}$ denote its eigenvectors. $\lambda_{1}$ is always 00 and $v_{1}$ is always $\vec{1}/\sqrt{n}$.
The Transformer architecture is as defined in Section 6, with $k=n$. We increase the dimension of $\Phi_{l}$ to $(2n)\times n$, and thus the dimension of $Z_{l}$ to $(d+n+n)\times n$. We read out the last $n$ rows of $Z_{L}$ as output. The purpose of increasing the dimension is to make the architecture identical to the one used in the experiments in Section 3.5; the construction in Lemmas 6 and Corollary 7 extend to this setting by setting appropriate parameters to 00. In addition, we also normalize ${\left[Z_{l}\right]}_{1...d}$ each layer by its Frobenius norm, i.e., ${\left[Z_{l}\right]}_{1...d}\leftarrow{\left[Z_{l}\right]}_{1...d}/\|{\left[Z_{l}\right]}_{1...d}\|_{F}$ (the proof of Lemma 6 still holds as subspace iteration is scaling-invariant). The input to the Transformer is $Z_{0}^{\top}=\begin{bmatrix}B,&\Phi_{0}\end{bmatrix}$, and we make $\Phi_{0}\in\mathbb{R}^{2n\times n}$ a trainable parameter along with $W^{V},W^{Q},W^{K},W^{R}$. We define $\text{loss}_{i}:=\mathbb{E}{\left[\min\left\{\|\phi_{L,i}-v_{i}\|_{2}^{2},\|\phi_{L,i}+v_{i}\|_{2}^{2}\right\}\right]}$, where $\phi_{L,i}$ is the $i^{th}$ row of $\Phi_{L}$ and expectation is taken with respect to randomness in sampling the graph; the $\min$ is required because eigenvectors are invariant to sign-flips. We train and evaluate the Transformer on two kinds of losses: $\text{loss}_{1-5}:=\frac{1}{5}\sum_{i=1}^{5}\text{loss}_{i}$ and $\text{loss}_{1-10}:=\frac{1}{10}\sum_{i=1}^{10}\text{loss}_{i}$. We plot the results in Figure 2, and summarize our findings below:
@@ 5 Parameter-Efficient Implementation
As explained in Section 3.1, the sizes of the $W^{Q},W^{K},W^{V},W^{R}$ matrices can scale with $O(n^{4})$ in the worst case, where $n$ is the number of nodes. This makes the memory requirement prohibitive. To alleviate this, we present below in (7) a more efficient Transformer implementation than the one defined in (3). The representation in (7) is strictly more constrained than (3), but it is still expressive enough to implement all the previous constructions. For each layer $l$, let $\alpha^{V}_{l}$, $\alpha^{Q}_{l}$, $\alpha^{K}_{l}$, $\alpha^{R}_{l}$ be scalar weight parameters, and let $W_{l}^{V,\Phi}$, $W_{l}^{Q,\Phi}$, $W_{l}^{K,\Phi}$, $W_{l}^{R,\Phi}\in\mathbb{R}^{2k\times 2k}$. Let $B_{l},\Phi_{l}$ evolve as
$B_{l+1}^{\top}=(1+\alpha^{R}_{l})B_{l}^{\top}+\alpha^{V}_{l}B_{l}^{\top}{\left(\alpha^{Q}_{l}\alpha^{K}_{l}B_{l}B_{l}^{\top}+\Phi_{l}{W_{l}^{Q,\Phi}}^{\top}W_{l}^{K,\Phi}\Phi_{l}^{\top}\right)}$ (7)
@@
with initialization $B_{0}:=B$, and $\Phi_{0}=[\Psi,0_{k\times n}]$. We verify that (7) matches (3), with $Z_{l}=\begin{bmatrix}B_{l}^{\top},\Phi_{l}^{\top}\end{bmatrix}$; the difference is that certain blocks of $W^{V}_{l}$, $W^{Q}_{l}$, $W^{K}_{l}$, $W^{R}_{l}$ constrained to be zero or some scaling of identity. Nonetheless, we verify that the constructions in Lemmas 1, 2, 3 and 6 can all be realized within the constrained dynamics (7). We prove this in Lemma 9 in Appendix 12. In Figure 15, we show that the efficient implementation (7) performs similarly (and in many cases better than) the standard implementation (3) on all the synthetic experiments.
Besides the reduced parameter size, we highlight two additional advantages of (7):
@@
    Let ${\sf TF}^{B}_{l}({\color[rgb]{0,0,1}B},{\color[rgb]{0,0,1}\Phi})$ (resp ${\sf TF}^{\Phi}_{l}({\color[rgb]{0,0,1}B},{\color[rgb]{0,0,1}\Phi}))$ be defined as the value of $B_{l}^{\top}$ (resp $\Phi_{l}^{\top}$) when initialized at $B_{0}={\color[rgb]{0,0,1}B}$ and $\Phi_{0}={\color[rgb]{0,0,1}\Phi}$ under the dynamics (7). Let $U\in\mathbb{R}^{d\times d}$ be some permutation matrix over edge indices. Then